# Training Large Neural Networks With Low-Dimensional Error Feedback

### v3

@@ 6.   Discussion
Our work demonstrates that neural networks can be trained effectively using minimal error signals constrained to the task’s intrinsic dimensionality rather than the higher dimensionality of the network’s representations. By adopting a factorized version of Feedback Alignment with low-rank matrices and training both left and right spaces of the feedback matrix, we showed that deep networks—linear, nonlinear, and convolutional—can match the performance of full backpropagation even under stringent error-dimensionality constraints. This finding highlights that the essential information required for learning is tied to the complexity of the task, measured by the number of output classes, and does not scale with the number of parameters or size of the model. Additionally, we revealed that constraining error dimensionality influences neural representations, providing a potential explanation for biological phenomena, such as center-surround receptive fields in the retina.
This work aims to explore potential mechanisms for implementing gradient descent in the brain. Recognizing that high-dimensional feedback is not necessary for effective learning is a significant step toward developing more flexible and biologically realistic models of learning. This insight suggests that the brain might utilize low-dimensional error signals to drive learning processes, aligning with anatomical and physiological constraints.
@@
Our work aims to reconcile gradient-based learning with the anatomical constraints of cortical circuits. In the cortex, feedback connections are generally sparser, less structured, and more selective than their feedforward counterparts [32], suggesting that learning may rely on a restricted top-down error signal. Several alternative frameworks have been proposed to explain learning in hierarchical biological circuits. Predictive coding models suggest that the cortex continuously generates predictions about sensory input and computes discrepancies, or prediction errors, between expected and actual signals [33, 34] . Equilibrium propagation [35] offers another biologically inspired alternative, demonstrating how energy-based models can approximate backpropagation through local weight updates. While these approaches provide valuable insights into neural computation, they deviate from the conventional framework of gradient-based optimization. In contrast, our approach remains firmly within the gradient descent paradigm, allowing direct loss minimization. This key property enables its seamless application across diverse architectures, from fully connected nonlinear networks to convolutional models and transformer-based attention mechanisms.
Between reward-based learning and full gradient descent.
By showing that deep networks can be trained with low-dimensional error signals, our framework bridges the gap between gradient-based learning and reward-based learning. Studies have suggested that dopaminergic neurons in the ventral tegmental area (VTA) encode reward signals that capture multiple task attributes rather than a single scalar value [36, 37, 38]. Our findings provide a theoretical foundation for moving beyond scalar reward-based learning toward a framework where low-dimensional reward signals carry richer information about the underlying loss (or gain). This perspective positions our model as an intermediary between reinforcement learning’s scalar feedback and the full backpropagation of error gradients. More broadly, our results suggest that learning could emerge through the strategic compression of error signals into task-relevant representations, offering a biologically plausible alternative to high-dimensional gradient propagation.
Aligning to the error space.
On its surface, our novel learning rule is similar to previous Feedback Alignment (FA) schemes. However, it offers a conceptual novelty. In traditional FA, learning the feedback weights aims to align them with the feedforward weights [5, 10], mirroring full backpropagation. In contrast, by factorizing the feedback matrix $B=QP$, we also align the row space of the feedback matrix with the source of the error, thereby improving the quality of the error signal itself.
Computational complexity.
Despite the additional requirement of training the feedback matrices, the computational cost of using low-rank feedback does not increase significantly. The updates for the low-rank matrices are linear in the number of neurons or filters per layer, while the forward pass and the primary weight updates remain quadratic [39]. This efficiency makes our approach practical for large-scale networks and suggests that similar mechanisms could be feasible in biological neural networks.
Gradient descent in high dimensions.
Our findings invite a rethinking of gradient descent dynamics in large, overparameterized networks. Typically, the weight dynamics during training are high-dimensional [40]. However, when the error signal is low-dimensional, the weight updates in each layer are confined to a much lower-dimensional subspace. This constraint could have implications for understanding the generalization capabilities of neural networks, as it suggests that effective learning does not require exploring the full parameter space. Indeed, previous studies found that the Hessian of the loss function shows low-rank structure [41]. Exploring this connection could open new avenues for understanding the dynamics of gradient descent in high-dimensional loss spaces [42].
Exploring the potential of low-dimensional feedback in more complex architectures.
This approach not only improves learning efficiency but may also act as a form of regularization, as can be seen from the improved performance of low-rank feedback matrices on visual transformers—the possibility of using restricted error signals as a form of regularization warrants further investigation.
In summary, our findings not only challenge the conventional reliance on high-dimensional error propagation but also pave the way for the development of biologically inspired, computationally efficient learning algorithms that hold promise for advancing both artificial intelligence and our understanding of brain function." This would leave the reader with a clear impression of both the scientific and translational impact of your work
@@ Acknowledgments
This work was supported by the Gatsby Charitable Foundation and the Azrieli Foundation.
@@ Appendix A Full linear theory
in this section, we provide the analysis of the linear networks for both Feedback Alignment and LDFA
@@
Full derivation of low-dimensional feedback alignment (LDFA) learning dynamics
@@
$\tau\frac{d\bar{Q}}{dt}$ $=\bar{W}_{1}\left(S-\bar{W}_{2}\bar{W}_{1}\right)^{\top}\bar{P}^{\top}$
@@ Appendix B Learning rule implementation
Our implementation and optimization were conducted entirely using PyTorch. For the LDFA layers, we initialized all weights using Kaiming uniform initialization. We modified the backward pass by adjusting the gradients with respect to the input, ensuring they align with our proposed update rule. In the output layer, we learn the projection matrix $\boldsymbol{P}$ to capture the principal directions of the target labels $\boldsymbol{y}$. In the hidden layers, $\boldsymbol{P}$ is learned to project onto the principal directions of the error signal from the subsequent layer, represented as $\boldsymbol{\delta}_{l+1}$. Specifically, for a layer $l$ with input dimension $n$, output dimension $m$, and rank constraint $r$, we proceed as follows:
Input: Error signal $\boldsymbol{\delta}_{l+1}\in\mathbb{R}^{b\times m}$, activations $\boldsymbol{h}_{l}\in\mathbb{R}^{b\times n}$, matrices $Q_{l}\in\mathbb{R}^{n\times r}$, $P_{l}\in\mathbb{R}^{r\times m}$, weight matrix $W_{l}\in\mathbb{R}^{m\times n}$
@@
Compute covariance matrix:
@@
Compute ’gradient’ w.r.t. input:
[…]
Compute updates for $Q_{l}$ and $P_{l}$:
@@
Compute update for $W_{l}$:
Algorithm 1 Modified Backward Pass for LDFA Layer
The derivation of the dLDFA (directed restricted adaptive feedback) follows the same procedure, with the key difference being that the error signal $\boldsymbol{\delta}$ does not necessarily originate from the next layer; instead, it can come from any subsequent layer.
@@ Appendix C Numerical experiments
In Figure [3.a], we present the results of training a network with four hidden layers, each containing 512 neurons, and an output layer with 10 neurons on the CIFAR-10 dataset. The network was trained using LDFA, as described earlier, without rank constraints, except for one layer at a time. For comparison, we also trained the network using standard backpropagation as a baseline. All networks were trained with a batch size of 32, a learning rate of $6\times 10^{-4}$, and weight decay of $4\times 10^{-4}$. The Adam optimizer with AMSGrad was used, with training conducted for 160 epochs and an exponential learning rate decay factor of 0.975. Each experiment was repeated 10 times.
@@
For the results shown in Figure [3.b], we trained the same network with four hidden layers, each containing 512 neurons. This time, we applied rank constraints to all layers simultaneously, with rank values $r=64,32,16,10$.
To further demonstrate that the network still utilizes high-rank representations, we also trained a variant of the network with 64 neurons in each hidden layer, without applying any rank constraints. All training was conducted with a batch size of 32, a learning rate of $6\times 10^{-4}$, and weight decay of $4\times 10^{-4}$. The Adam optimizer with AMSGrad was used, with training carried out for 160 epochs and an exponential learning rate decay factor of 0.975. Each experiment was repeated 10 times.
CIFAR-100 Sub-sampling
For the results shown in Figure [3.c], we trained the same model on the CIFAR-100 dataset, sampling different numbers of classes $d$, with $d=50,75,100$. For each sub-sample, we trained the model while applying rank constraints to all layers, using various rank values. This was done to demonstrate that the dimensionality of the error signal depends on the task dimensionality $d$. All optimizations were performed with a batch size of 32, a learning rate of $6\times 10^{-4}$, and weight decay of $4\times 10^{-4}$. The Adam optimizer with AMSGrad was used, with training conducted for 160 epochs and an exponential learning rate decay factor of 0.975. Each training run was repeated 5 times.
Direct LDFA
For the results shown in Figure [3.d], we trained the same model using dLDFA. In one experiment, we propagated the error signal from the last layer to all preceding layers. In a separate experiment, we propagated the error signal directly from the penultimate layer to all earlier layers, applying rank constraints to these layers while keeping the last layer at full rank. All optimizations were performed with a batch size of 32, a learning rate of $6\times 10^{-4}$, and weight decay of $4\times 10^{-4}$. The Adam optimizer with AMSGrad was used, training for 160 epochs with an exponential learning rate decay factor of 0.975. Each training run was repeated 5 times
Convolutional Neural Networks
To extend our LDFA algorithm to convolutional layers, we apply the rank constraint to the number of channels in the error signal. This effectively constrains the dimensionality of the error signal across the spatial dimensions. Similar to the implementation for fully connected layers in Algorithm 1, we modify the backward pass in the same way for convolutional layers.
@@
The key difference in the convolutional context is how the projection matrix $\boldsymbol{P}$ operates on the error signals. In convolutional layers, $\boldsymbol{P}$ functions as a $1\times 1$ convolutional filter, projecting the error signal at each spatial location from $m$ channels down to $r$ channels. This reduces the error signal’s dimensionality to $r$ per pixel, adhering to the rank constraint.
As in the fully connected case, the matrix $\boldsymbol{P}$ is learned using Oja’s rule. However, the covariance matrix $\boldsymbol{C}$ is computed over both the batch and spatial dimensions—that is, across all pixels in all images within the batch. This approach captures the covariance structure of pixel representations more effectively, enabling $\boldsymbol{P}$ to project the error signals appropriately in the convolutional setting.
VGG-like Architecture
We trained a VGG-like convolutional neural network comprising four convolutional blocks, as detailed in Table C1. For the experiments presented in Figure 4(a), a rank constraint was applied exclusively to the final convolutional block, which consists of 512 channels. In contrast, for the results shown in Figure 4(b), rank constraints were imposed on all layers, reducing the rank to 1/2, 1/4, and 1/8 of the original channel dimensions.
@@
Each network was trained using the Adam optimizer with a learning rate of $5\times 10^{-4}$, a weight decay of $5\times 10^{-5}$, and an exponential learning rate decay factor of 0.98. Training was conducted for 250 epochs, with each experiment repeated five times to ensure robustness.
Vision Transformer (ViT)
We implemented a Vision Transformer (ViT) model that consists of seven self-attention blocks, each with an embedding dimension of 384 and four attention heads. The input images were partitioned into nonoverlapping patches, which were linearly projected into a 384-dimensional feature space. These patch embeddings were augmented with learned positional encodings and processed through a stack of seven transformer encoder layers. Each encoder block employed multi-head self-attention with four heads, followed by pre-norm layer normalization and dropout for regularization. The final representation was obtained from a learned class token, which was subsequently processed by a fully connected layer for classification.
Training was carried out using a batch size of 32, a learning rate of $3\times 10^{-4}$, and a weight loss of $1\times 10^{-4}$. We used the Adam optimizer with AMSGrad, training for 150 epochs with an exponential learning rate decay factor of 0.98. Each experiment was repeated five times to ensure robustness.
@@
Table C1: CNN Architecture
Layer(s) Output Size Details
Input $32\times 32\times 3$
Convolutional Block 1
Conv2D + ReLU $32\times 32\times 64$ $3\times 3$ conv, 64 filters, padding=1
BatchNorm2D $32\times 32\times 64$
Conv2D + ReLU $32\times 32\times 64$ $3\times 3$ conv, 64 filters, padding=1
BatchNorm2D $32\times 32\times 64$
MaxPool2D $16\times 16\times 64$ $2\times 2$ max pool, stride=2
Convolutional Block 2
Conv2D + ReLU $16\times 16\times 128$ $3\times 3$ conv, 128 filters, padding=1
BatchNorm2D $16\times 16\times 128$
Conv2D + ReLU $16\times 16\times 128$ $3\times 3$ conv, 128 filters, padding=1
BatchNorm2D $16\times 16\times 128$
MaxPool2D $8\times 8\times 128$ $2\times 2$ max pool, stride=2
Convolutional Block 3
Conv2D + ReLU $8\times 8\times 256$ $3\times 3$ conv, 256 filters, padding=1
BatchNorm2D $8\times 8\times 256$
Conv2D + ReLU $8\times 8\times 256$ $3\times 3$ conv, 256 filters, padding=1
BatchNorm2D $8\times 8\times 256$
MaxPool2D $4\times 4\times 256$ $2\times 2$ max pool, stride=2
Convolutional Block 4
Conv2D + ReLU $4\times 4\times 512$ $3\times 3$ conv, 512 filters, padding=1
BatchNorm2D $4\times 4\times 512$
Conv2D + ReLU $4\times 4\times 512$ $3\times 3$ conv, 512 filters, padding=1
BatchNorm2D $4\times 4\times 512$
AdaptiveAvgPool2D $1\times 1\times 512$ Output size $(1,1)$
Flatten $512$ Flatten to vector
Classifier
Fully Connected + ReLU $256$ Linear layer, $512\rightarrow 256$
Dropout $256$ Dropout probability $p=0.4$
Fully Connected $C$ Linear layer, $256\rightarrow C$

### v4

@@ 7.  Discussion
Our central result is that deep networks do not require transmitting the full layerwise gradient: it is sufficient to deliver a low-dimensional teaching signal to each layer, provided the error is projected into task-relevant directions. In a solvable linear theory, fixed low-rank feedback generically makes the stationarity conditions underdetermined and introduces spurious fixed points. Learning the feedback factors aligns the transmitted subspace with the evolving error and restores backpropagation-like convergence across nonlinear networks, including CNNs and vision transformers.
This perspective cleanly separates a normative question (which error subspace should be communicated) from an implementational one (how that subspace can be learned). We first derive feedback learning as a low-rank factorization objective whose updates do not scale with batch size, and then show how essentially the same subspace can be learned with synapse-local plasticity driven by the arriving error, enabling direct error projection through non-mirrored pathways and revealing error dimensionality as an inductive bias that can shape early receptive fields.
Computational efficiency.
Because low-rank feedback constrains the error channel itself, the expensive multiplication by $W_{l}^{\top}$ in the backward pass is replaced by two thin multiplies, reducing per-example cost from $O(n_{l}n_{l+1})$ to $O(r(n_{l}+n_{l+1}))$ and making backward FLOPs scale linearly with the effective error dimensionality $r$. End-to-end, this creates a trade-off in which decreasing the rank $r$ lowers computational cost, but excessive compression degrades learning efficiency, requiring more samples. Consequently, total compute exhibits a U-shaped dependence on $r$, with an intermediate rank minimizing FLOPs-to-accuracy in our transformer experiments. This positions low-dimensional feedback as a structured alternative to gradient sparsification and compression methods such as meProp [66] and Deep Gradient Compression [39]. Rather than sparsifying a full gradient after it is formed, we learn the projection that determines which error directions are represented upstream, so the savings come from never materializing or transporting the discarded components.
Aligning to the error space.
At first glance, our local rule resembles adaptive variants of Feedback Alignment (FA), where the feedback weights are trained to better match the feedforward weights and thereby approximate backpropagation [38, 1]. The factorization $B_{l}=Q_{l}P_{l}$ reveals a different target: $Q_{l}$ only mixes an $r$-dimensional teaching channel into layer $l$, whereas $P_{l}$ determines which output-error directions can be represented and transmitted upstream (the subspace spanned by $P_{l}^{\top}$). Learning $P_{l}$ therefore aligns the transmitted error subspace with the dominant modes of the current residual, improving the information content of the teaching signal itself. This perspective explains why fixed low-rank feedback can induce spurious fixed points, and why subspace learning resolves them: the bottleneck fails when it discards task-relevant error directions, not merely when it misaligns with $W_{l}^{\top}$.
Low-rank adaptation (LoRA) in large language models.
Recent parameter-efficient fine-tuning methods for large language models, most notably LoRA, achieve strong adaptation by freezing pretrained weights and learning low-rank updates $\Delta W$ within transformer layers [26], with practical extensions that push memory efficiency further (e.g., quantization with QLoRA) and adaptive rank allocation) [17, 41]. Our approach is complementary in where the low-rank constraint is imposed. Here, instead of restricting the forward update parameterization while retaining full backpropagated errors, we restrict and learn the error subspace communicated to each layer. This provides a middle ground in which the number of trainable forward parameters can grow (increasing representational capacity) while the cost of credit assignment remains controlled by the feedback rank.
@@
Our work addresses a basic anatomical mismatch between backpropagation and cortex: feedback pathways are not organized as a synapse-wise transpose of the forward weights and need not deliver a neuron-by-neuron teaching signal [43]. Predictive coding and equilibrium propagation propose mechanistic routes by which local plasticity, together with inference dynamics, can approximate gradient-based learning in hierarchical circuits [53, 69, 61]. Our contribution is complementary: we remain in the standard objective-minimization setting, but treat the teaching pathway as a constrained resource and learn a low-dimensional projection of error signals that preserves task-relevant directions. This yields effective credit assignment under severe bandwidth limits and, via direct error projection, remains compatible with indirect and non-reciprocal delivery of instructive signals.
Low-dimensional teaching signals and reward pathways.
The same perspective offers a middle ground between scalar reinforcement signals and full gradient transport. Midbrain dopamine has long been linked to reward-prediction-error teaching signals, yet recordings increasingly suggest heterogeneous coding in which dopaminergic populations and terminals reflect multiple sensory, motor, and cognitive variables beyond a single scalar reward [13, 49, 19]. In our framework, such signals can be interpreted as a low-dimensional teaching vector: a compressed, broadcast error representation whose efficacy depends on learning the appropriate projection into each layer’s plastic synapses. This suggests a concrete computational role for low-dimensional neuromodulatory feedback, without requiring the brain to propagate a full high-dimensional gradient through deep circuits.
Feedback structure as an inductive bias.
Inductive biases in deep learning are usually discussed as properties of the forward architecture, where weight sharing, locality, and symmetry constraints bias learning toward particular solution classes and improve generalization [71, 8, 64, 23]. Our results suggest that the teaching channel can play a similar role by constraining and shaping the feedback pathway. The feedback path controls which error directions are expressible upstream, thereby biasing optimization toward representations that are easiest to learn under the available error bandwidth. This viewpoint is consistent with recent analyses showing that feedback-alignment-style learning exhibits implicit regularization and can select among degenerate low-loss solutions [54]. In our setting, the receptive-field experiments provide a concrete illustration that modifying only the feedback pathway can systematically alter learned symmetries. A promising direction is therefore to treat feedback design as a new knob for algorithmic inductive bias: for example, imposing topographic, convolutional, modular, or sign-constrained structure on low-rank feedback may preferentially promote invariances, robustness, or task-specific feature reuse, while maintaining computational efficiency through a controlled error bottleneck.
In summary, we argue that the essential constraint for scalable credit assignment is not access to a full, layerwise gradient but access to a learned low-dimensional teaching channel that preserves task-relevant error directions. This reframes feedback as a resource that can be optimized, offering a principled route to more efficient training when backward computation or communication is limiting, and a concrete hypothesis for how anatomically restricted pathways could still support deep learning. More broadly, our results invite future work that treats the structure of feedback, not only the forward architecture, as a design axis that shapes both efficiency and the inductive biases of learned representations.
@@ Acknowledgments
This work was supported by the Gatsby Charitable Foundation.
@@ Appendix A Full linear theory
In this section, we provide the analysis of the linear networks for both Feedback Alignment and LDFA.
@@
Derivation of Normative Low-Rank Feedback Learning
Here we derive the learning dynamics for the normative variant of LDFA, where the feedback matrix $B=QP$ is trained to explicitly approximate the transpose of the forward weights $W_{2}^{\top}$. The feedback factors $Q\in\mathbb{R}^{k\times r}$ and $P\in\mathbb{R}^{r\times m}$ are updated to minimize the reconstruction loss:
$\mathcal{L}_{B}=\frac{1}{2}\|QP-W_{2}^{\top}\|_{F}^{2}.$ (A27)
The gradients with respect to the feedback factors are:
$\frac{\partial\mathcal{L}_{B}}{\partial P}$ $=Q^{\top}(QP-W_{2}^{\top}),$ (A28)
$\frac{\partial\mathcal{L}_{B}}{\partial Q}$ $=(QP-W_{2}^{\top})P^{\top}.$
Assuming gradient descent with a time constant $\tau_{B}$, the continuous-time dynamics are:
$\tau_{B}\frac{dP}{dt}$ $=Q^{\top}(W_{2}^{\top}-QP),$ (A29)
$\tau_{B}\frac{dQ}{dt}$ $=(W_{2}^{\top}-QP)P^{\top}.$
We now rotate these dynamics into the singular basis of $\Sigma_{io}$ using the same transformations as above: $W_{2}=U\bar{W}_{2}$ (implying $W_{2}^{\top}=\bar{W}_{2}^{\top}U^{\top}$) and $P=\bar{P}U^{\top}$. Note that $Q$ operates in the hidden space and is not rotated by $U$.
Substituting these into the dynamics for $P$:
$\tau_{B}\frac{d(\bar{P}U^{\top})}{dt}$ $=Q^{\top}\left(\bar{W}_{2}^{\top}U^{\top}-Q\bar{P}U^{\top}\right)$ (A30)
$=Q^{\top}\left(\bar{W}_{2}^{\top}-Q\bar{P}\right)U^{\top}.$ (A31)
Multiplying by $U$ on the right yields the rotated dynamics for $\bar{P}$:
$\tau_{B}\frac{d\bar{P}}{dt}=Q^{\top}\left(\bar{W}_{2}^{\top}-Q\bar{P}\right).$ (A32)
Similarly for $Q$:
$\tau_{B}\frac{dQ}{dt}$ $=\left(\bar{W}_{2}^{\top}U^{\top}-Q\bar{P}U^{\top}\right)(\bar{P}U^{\top})^{\top}$ (A33)
$=\left(\bar{W}_{2}^{\top}-Q\bar{P}\right)U^{\top}U\bar{P}^{\top}$ (A34)
$=\left(\bar{W}_{2}^{\top}-Q\bar{P}\right)\bar{P}^{\top}.$ (A35)
Together with the forward weight updates derived in Eq. (A19) and (A25), this forms the closed dynamical system described in Eq. (10) of the main text.
Full derivation of local low-dimensional feedback alignment (LDFA) learning dynamics
@@
$\tau\frac{dQ}{dt}$ $=\bar{W}_{1}\left(S-\bar{W}_{2}\bar{W}_{1}\right)^{\top}\bar{P}^{\top}$
@@ Appendix B Learning rule implementation
Our implementation and optimization were conducted entirely using PyTorch. For the LDFA layers, we initialized all weights using Kaiming uniform initialization. We modified the backward pass by adjusting the gradients with respect to the input, ensuring they align with our proposed update rule as well as calculating the gradients of $P$ and $Q$.
We implemented two variants of the algorithm. In the **Normative** version (Algorithm 1), $P$ and $Q$ are updated to explicitly minimize the reconstruction error $\|QP-W^{\top}\|$, ensuring the feedback pathway approximates the transpose of the forward weights. In the **Local** version (Algorithm 2), $P$ and $Q$ are updated using biologically plausible, local plasticity rules (Oja’s rule) that align the feedback subspace with the principal components of the error signal.
Input: Error signal $\boldsymbol{\delta}_{l+1}\in\mathbb{R}^{b\times m}$, activations $\boldsymbol{h}_{l}\in\mathbb{R}^{b\times n}$, matrices $Q_{l}\in\mathbb{R}^{n\times r}$,
     $P_{l}\in\mathbb{R}^{r\times m}$, weight $W_{l}\in\mathbb{R}^{m\times n}$
@@
Compute ’gradient’ w.r.t. input:
$\text{grad\_input}\leftarrow\left(Q_{l}P_{l}\boldsymbol{\delta}_{l+1}^{\top}\right)^{\top}\in\mathbb{R}^{b\times n}$;
Compute update for $W_{l}$:
Compute updates for $Q_{l}$ and $P_{l}$ (Reconstruction):
$E\leftarrow W_{l}^{\top}-Q_{l}P_{l}$;
$\Delta Q_{l}\leftarrow EP_{l}^{\top}$;
$\Delta P_{l}\leftarrow Q_{l}^{\top}E$;
Algorithm 1 Backward Pass for Normative LDFA (Approximating $W^{\top}$)
In the local version (Algorithm 2), we learn the projection matrix $\boldsymbol{P}$ to capture the principal directions of the target labels $\boldsymbol{y}$ (in the output layer) or the error signal $\boldsymbol{\delta}_{l+1}$ (in hidden layers). Specifically, for a layer $l$ with input dimension $n$, output dimension $m$, and rank constraint $r$, we proceed as follows:
Input: Error signal $\boldsymbol{\delta}_{l+1}\in\mathbb{R}^{b\times m}$, activations $\boldsymbol{h}_{l}\in\mathbb{R}^{b\times n}$, matrices $Q_{l}\in\mathbb{R}^{n\times r}$,
     $P_{l}\in\mathbb{R}^{r\times m}$, weight $W_{l}\in\mathbb{R}^{m\times n}$
Compute ’gradient’ w.r.t. input:
$\text{grad\_input}\leftarrow\left(Q_{l}P_{l}\boldsymbol{\delta}_{l+1}^{\top}\right)^{\top}\in\mathbb{R}^{b\times n}$;
Compute update for $W_{l}$:
Compute covariance matrix for P update:
@@
Compute updates for $Q_{l}$ and $P_{l}$ (Local):
@@
Algorithm 2 Backward Pass for Local LDFA (Oja’s Rule)
The derivation of the dLDFA (directed restricted adaptive feedback) follows the same procedure as Algorithm 2, with the key difference being that the error signal $\boldsymbol{\delta}$ does not necessarily originate from the next layer; instead, it can come from any subsequent layer.
@@ Appendix C Numerical experiments
C.1 Common Architectures and Training Details
Across our experiments, we utilized three primary architectures: a fully connected Multi-Layer Perceptron (MLP), a VGG-style Convolutional Neural Network (CNN), and a Vision Transformer (ViT). Unless otherwise stated in the specific experiment subsections, the architectures follow the specifications below.
Multi-Layer Perceptron (MLP)
For the dense network experiments on CIFAR-10 and CIFAR-100, we used a network with 4 hidden layers.
  • Structure: Input $\to$ [Linear(512) $\to$ ReLU] $\times$ 3 $\to$ Linear(Output).
  • Hidden Width: 512 neurons (unless specified otherwise for scaling experiments).
  • Initialization: Kaiming Uniform.
Convolutional Neural Network (VGG-Like)
We trained a VGG-style network with four convolutional blocks and batch normalization. The detailed architecture is provided in Table C1.
Table C1: VGG-style CNN Architecture used for CIFAR-10 experiments.
Block Layer Type Details
Block 1 $2\times$ (Conv2d + BN + ReLU) $3\times 3$ kernel, 64 filters, padding=1
MaxPool2d $2\times 2$ kernel, stride=2
Output Size $16\times 16\times 64$
Block 2 $2\times$ (Conv2d + BN + ReLU) $3\times 3$ kernel, 128 filters, padding=1
MaxPool2d $2\times 2$ kernel, stride=2
Output Size $8\times 8\times 128$
Block 3 $2\times$ (Conv2d + BN + ReLU) $3\times 3$ kernel, 256 filters, padding=1
MaxPool2d $2\times 2$ kernel, stride=2
Output Size $4\times 4\times 256$
Block 4 $2\times$ (Conv2d + BN + ReLU) $3\times 3$ kernel, 512 filters, padding=1
AdaptiveAvgPool2d Output size $(1,1)$
Flatten Vector size 512
Classifier Linear + ReLU $512\to 256$
Dropout $p=0.4$
Linear $256\to\text{Num Classes}$
Vision Transformer (ViT)
We utilized the Vision Transformer implementation from the timm library. The default configuration used across experiments (unless explicitly modified for scaling tests) is detailed in Table C2.
Table C2: Vision Transformer (ViT) hyperparameters.
Parameter Value
Patch Size 4
Input Channels 3
Embedding Dimension 384 (variable in scaling experiments)
Depth 8 layers
Number of Heads 8
MLP Ratio 2.0
QKV Bias True
Drop Rate 0.1
Attention Drop Rate 0.1
Drop Path Rate 0.1
C.2 Experiments with Normative Feedback
In this section, we detail the experiments corresponding to Section 4 of the main text, where the feedback matrices $B_{l}=Q_{l}P_{l}$ are trained to minimize the reconstruction error $\|Q_{l}P_{l}-W_{l}^{\top}\|_{F}^{2}$ (Algorithm 1).
Dual Optimization Strategy
In all normative experiments, we employed two separate optimizers: one for the forward weights ($W_{l}$) and one for the backward feedback factors ($Q_{l},P_{l}$). Both used the AdamW algorithm but were configured with different learning rates and schedules to ensure the feedback pathway effectively tracked the evolving forward weights.
Convolutional Networks (Layer-wise Constraints)
To evaluate the impact of error compression on deep convolutional networks (Fig. 3a), we trained the VGG-like architecture defined in Table C1 on the CIFAR-10 dataset.
  • Rank Constraints: We replaced the standard backward pass in every convolutional layer with a low-rank feedback channel. The rank $r$ was set to a fixed fraction of the layer’s channel width $n$ (i.e., $r\in\{n/2,n/4,n/8\}$).
  • Feedback Kernel: For convolutional layers, the projection $P_{l}$ operates as a $1\times 1$ convolution, reducing the error signal from $m$ channels to $r$ channels per spatial location.
Vision Transformers (ViT)
We trained the Vision Transformer using the dual-optimizer setup described below. The training duration was set to 250 epochs for CIFAR-10 and 300 epochs for CIFAR-100.
  • Scope: Low-rank feedback was applied to all learnable linear transformations, including Attention (Q, K, V) and MLP layers.
  • Forward Optimization: The forward weights were optimized using AdamW with a base learning rate of $3\times 10^{-4}$ and weight decay of $0.1$.
  • Backward Optimization: The feedback factors were optimized using AdamW with a base learning rate of $0.015$ and no weight decay.
  • Scheduling: Both optimizers utilized a warmup phase of 10 epochs with a start factor of 0.1 (via linear scheduling), followed by a cosine annealing schedule decaying to a minimum learning rate of $1\times 10^{-9}$.
  • CIFAR-100 Adaptation: For CIFAR-100 experiments, the backward optimizer’s base learning rate was adjusted to $0.005$.
C.3 Experiments with Local Feedback
In this section, we verify the efficacy of the local learning rule (Algorithm 2), where the feedback projection $\boldsymbol{P}_{l}$ is learned via an error-driven Oja’s rule to capture the principal subspace of the teaching signal.
Multi-Layer Perceptron (MLP) Experiments
Experimental Setup
Unless otherwise noted, we used the standard MLP architecture described in the Common Architectures section (4 hidden layers, 512 neurons). All models were trained for 160 epochs using the Adam optimizer with AMSGrad, a batch size of 32, a learning rate of $6\times 10^{-4}$, and a weight decay of $4\times 10^{-4}$. The learning rate followed an exponential decay with a factor of 0.975.
@@ C.3 Experiments with Local Feedback
To analyze the sensitivity of different depths to error compression, we applied the LDFA rank constraint to only one layer at a time, leaving all other layers unconstrained. We compared these distinct training runs against a baseline trained with standard backpropagation. Each experiment was repeated 10 times.
@@
We evaluated the performance when the entire error pathway is compressed by applying rank constraints simultaneously to all layers with ranks $r\in\{64,32,16,10\}$. To verify that the network relies on high-dimensional feedforward representations even when feedback is low-rank, we trained a control model where the feedforward width was physically reduced to 64 neurons (matching the feedback bottleneck) without rank constraints. These experiments were repeated 10 times.
Task Dimensionality (CIFAR-100)
To test the hypothesis that feedback rank scales with task dimensionality $d$, we trained the model on CIFAR-100 subsets with $d\in\{50,75,100\}$ classes. For each subset, we applied rank constraints to all layers. Each run was repeated 5 times.
Direct LDFA (dLDFA)
We evaluated the local rule’s ability to support broadcast error signals in two configurations:
  • Output Broadcast: The error signal $\boldsymbol{\delta}_{L}$ was propagated directly from the output layer to all preceding hidden layers.
  • Penultimate Broadcast: The error signal was propagated from the penultimate layer to all earlier layers (applying rank constraints), while the final layer was updated normally.
These runs were repeated 5 times.
Convolutional Neural Networks (CNN)
Methodology
@@
To extend LDFA to convolutional layers, we treat the rank constraint as a bottleneck on the number of error channels. The projection matrix $\boldsymbol{P}$ functions as a $1\times 1$ convolutional filter, projecting the error signal at each spatial location from $m$ channels down to $r$ channels. Crucially, the covariance matrix $C$ for Oja’s rule (Algorithm 2) is computed by aggregating over both the batch and spatial dimensions (i.e., across all pixels in all images). This ensures $\boldsymbol{P}$ captures the covariance structure of the pixel-wise error representations.
VGG-like Experiments
@@
We trained the VGG-style architecture defined in Table C1. Training was conducted for 250 epochs using Adam with a learning rate of $5\times 10^{-4}$, weight decay of $5\times 10^{-5}$, and an exponential decay factor of 0.98. All experiments were repeated 5 times.